# MSCPT: Few-shot Whole Slide Image Classification with Multi-scale and Context-focused Prompt Tuning

### v1

@@ Abstract
Multiple instance learning (MIL) has become a standard paradigm for weakly supervised classification of whole slide images (WSI). However, this paradigm relies on the use of a large number of labelled WSIs for training. The lack of training data and the presence of rare diseases present significant challenges for these methods. Prompt tuning combined with the pre-trained Vision-Language models (VLMs) is an effective solution to the Few-shot Weakly Supervised WSI classification (FSWC) tasks. Nevertheless, applying prompt tuning methods designed for natural images to WSIs presents three significant challenges: 1) These methods fail to fully leverage the prior knowledge from the VLM’s text modality; 2) They overlook the essential multi-scale and contextual information in WSIs, leading to suboptimal results; and 3) They lack exploration of instance aggregation methods. To address these problems, we propose a Multi-Scale and Context-focused Prompt Tuning (MSCPT) method for FSWC tasks. Specifically, MSCPT employs the frozen large language model to generate pathological visual language prior knowledge at multi-scale, guiding hierarchical prompt tuning. Additionally, we design a graph prompt tuning module to learn essential contextual information within WSI, and finally, a non-parametric cross-guided instance aggregation module has been introduced to get the WSI-level features. Based on two VLMs, extensive experiments and visualizations on three datasets demonstrated the powerful performance of our MSCPT.
Index Terms:
@@ I Introduction
Developing automated analysis frameworks using Whole Slide Images (WSIs) is crucial in clinical practice [1, 2, 3, 4], as WSIs are widely regarded as the “gold standard” for cancer diagnosis, typing, staging, and prognosis analysis [5, 6]. Given the enormous size of WSIs (roughly 40,000 × 40,000), multiple instance learning (MIL) [7] has become the dominant method. As shown in Fig. 1 a, traditional MIL-based methods typically follow a four-step paradigm: patch cutting, feature extraction, feature aggregation, and classification. Most MIL-based methods are conducted under weak supervision at the bag level, as creating instance-level labels is quite labor-intensive [8, 9]. This weak supervision paradigm has led to a problem: a large number of WSIs are required to train an effective model [10, 11]. In clinical practice, patient privacy concerns, rare diseases, and difficulty preparing pathology slides make accumulating a large number of WSIs very challenging [12, 13].
Vision-Language models (VLMs) have shown excellent generalization ability to downstream tasks [14, 15, 16, 17, 18, 19, 20]. Recently, researchers have proposed specialized VLMs for analyzing pathological images, including MI-Zero [21], PLIP [22], and Conch [23]. These VLMs, extensively pre-trained on abundant image-text pairs, contain significant prior knowledge. If the prior knowledge of VLMs can be fully exploited with a few training samples, it can partially alleviate the data scarcity problem in WSI classification tasks. Therefore, we aim to explore a novel “data-efficient” method based on the VLMs to improve the model’s performance on the Few-shot Weakly Supervised WSI Classification (FSWC) [8] task.
@@
However, existing prompt tuning methods for natural images (such as CoOp [25], CoCoOp [24], and MetaPrompt [26]) are only effective for single images (i.e., patch-level). Since each WSI typically contains tens of thousands of patches, these methods are ineffective for WSI-level tasks. Also, studies indicate that the multi-scale information [27] and the contextual information [28] in WSIs play a significant role in cancer analysis, but those methods fail to capture this crucial information. Additionally, in training VLMs, the image-text pairs contain more than just information about the category. They also include more details about the image, such as contextual properties of the object [14] and descriptions of the cellular microenvironment [22, 23]. However, existing prompt tuning methods have primarily focused on image category information, without emphasizing a detailed image content analysis, which has left the full potential of pathological VLMs underexplored.
@@
To address the aforementioned issue, we propose Multi-Scale and Context-focused Prompt Tuning (MSCPT) for WSI classification in weakly supervised and few-shot scenarios. Our framework fully leverages the characteristic of VLM training with image-text pairs at dual magnification scales: 1) At low magnification, we provide the VLM with pathological visual descriptions at the tissue level (such as the infiltration between tumor tissue and other normal tissues); 2) At high magnification, pathological visual descriptions at the cellular level (such as cell morphology, nuclear changes, and the formation of various organelles) are provided to the VLM. These pathological visual descriptions at multi-scale can help VLM identify regions that are helpful for cancer analysis and achieve optimal results even with limited training samples.
@@
As illustrated in Fig. 1 c, the core idea behind developing MSCPT is to incorporate prior knowledge at the tissue and cellular scales into the WSI-level tasks. Specifically, we first use a frozen large language model (LLM) to generate multi-scale pathological visual descriptions, leveraging them as prior knowledge.
@@
Secondly, we design a Multi-scale Hierarchical Prompt Tuning (MHPT) module to combine pathological visual descriptions from multi-scale hierarchically to enhance prompt effectiveness. Inspired by Metaprompt [26], a dual-path asymmetric framework is adopted, asymmetrically freezing the image encoder and text encoder at different scales for prompt tuning. This asymmetric framework enables us to freeze half of the encoder to reduce the number of trainable parameters. Specifically, MHPT contains low-level and high-level prompts for both low and high-magnification visual descriptions, as well as global trainable prompts. The MHPT module employs the transformer layers in the text encoder to effectively learn the interactions among three distinct prompts.
@@
Furthermore, the Image-text Similarity-based Graph Prompt Tuning (ISGPT) module is introduced to extract contextual information. Precisely, we do not follow previous approaches [29, 30] of using patch positions or patch feature similarity to construct graph neural networks (GNNs). We propose to use the similarity between patches and pathological visual descriptions as the basis for building GNNs. We believe that using image-text pairs to build GNNs is more effective for capturing global features than methods relying on patch positions and image feature similarity, and corresponding ablation experiments confirm this hypothesis.
@@
Finally, impressed by the powerful zero-shot capabilities of VLMs [22, 23, 21], we fully leverage the similarity between patches and pathological visual descriptions to aggregate instances. The Non-Parametric Cross-Guided Pooling (NPCGP) module, utilizing the Top-K algorithm for instance aggregation, is introduced to further reduce the risk of overfitting in few-shot scenarios. Overall, our contributions are summarised as follows:
@@
    MSCPT demonstrates that high-level concepts from pathological descriptions combined with low-level image representations can enhance few-shot weakly supervised WSI classification.
@@
    MSCPT achieves excellent performance by introducing only a limited number of trainable parameters ($\sim$0.9of the pre-trained VLM). Additionally, MSCPT is applicable to fine-tune any VLMs for WSI-level tasks.
@@
    Extensive experiments and visualizations on three datasets and two VLMs have confirmed that MSCPT’s performance is state-of-the-art in few-shot scenarios, surpassing other traditional MIL-based and prompt tuning methods.
@@
Figure 2: We develop MSCPT based on the dual-path asymmetric framework, which inputs patches and pathological visual descriptions from multi-scale to different encoders. MSCPT utilizes a large language model to generate multi-scale pathological visual descriptions. These descriptions are combined using Multi-scale Hierarchical Prompt Tuning (MHPT) to integrate information across multiple scales. Then Image-text similarity-based Graph Prompt Tuning (ISGPT) is employed to learn context information at each scale. Finally, Non-Parametric Cross-Guided Pooling (NPCGP) aggregates instances guided by pathological visual descriptions to achieve the final Whole Slide Image classification result.
II Related Work
II-A Multiple Instance Learning in Whole Slide Images
@@
Due to the high resolution of Whole Slide Images (WSIs) and the challenges of detailed labelling, weakly supervised methods based on Multiple Instance Learning (MIL) have emerged as the mainstream for WSI analysis. The MIL-based methods treat a WSI as a bag and all patches as instances, considering a bag positive if it contains at least one positive instance. Within the MIL framework, an aggregation step is required to aggregate all instances into bag features. The most primitive aggregation methods are non-parametric mean pooling and max pooling. However, since disease-related instances are a small fraction [31], those non-parametric aggregation methods treated all instances equally, causing useful information to be overwhelmed by irrelevant data. Subsequently, some attention-based methods (such as ABMIL [32], DSMIL [33] and CLAM [31]) were introduced, assigning different weights to each instance and aggregating them based on the weights. Furthermore, MIL methods based on Graph Neural Networks (GNNs) [29, 30] and Transformers [1, 34] had also been proposed to capture both local and global contextual information of WSIs. Those methods have shown significant improvements in recent years. Still, the cost of enhancing model performance is the increase in parameters, requiring a large amount of data to train a well-performing model. In many cases, training data faces a scarcity issue. Therefore, this paper proposes MSCPT, which leverages Vision-Language models combined with pathological descriptions from Large language models to enhance the performance in few-shot scenarios.
II-B Vision-Language Models
@@
Vision-Language models (VLMs) are rapidly developing in various fields. During training, VLMs use contrastive learning to reduce distances between paired image-text pairs and increase distances between unpaired ones. CLIP [14] collected over 400M image-text pairs from the internet and used contrastive learning to align them, resulting in compatibility across various tasks. Compared to natural images, gathering pairs of pathological images and corresponding descriptions is challenging. To address this issue, MI-Zero [21] first pretrains image and text encoders using unpaired data, and then aligns them in a common space using 33,480 pairs of pathological image-text pairs. Huang et al. gathered over 450K pathological image-text pairs from Twitter and LAION [35] and developed PLIP [22]. Lu et al. trained Conch [23] on over 1.17M pathological image-text pairs, and it performs well on downstream tasks. Pretrained VLMs have significant potential, but effective methods to leverage them for WSI-level tasks are lacking. In this paper, we propose using pathological visual descriptions as prior knowledge to unleash the potential of VLMs.
II-C Prompt Tuning in Vision-Language Models
@@
Prompt tuning has demonstrated remarkable efficiency and effectiveness, whether in text or multimodal [25, 24, 18]. CLIP demonstrated remarkable zero-shot performance with hand-crafted prompts, but the results can vary significantly depending on the prompt used due to their sensitivity to changes. Therefore, CoOp [25] and CoCoOp [24] proposed that the model itself should determine the choice of prompts. Khattak et al. argued that optimizing prompt tuning within a single branch is not ideal. They introduced MaPLe [18] as a solution to enhance the alignment between visual and language representations. Regrettably, these innovative methods are highly applicable to natural images but do not consider the enormous size of WSIs and the crucial multi-scale and contextual information needed for WSI analysis.
To our knowledge, Qu et al. have conducted research TOP [8] on the fine-tuning of CLIP for FSWC tasks. Shi et al. also proposed ViLa-MIL [36] based on CLIP, which helps with WSI classification by introducing multi-scale language prior knowledge. These two studies are exceptional, pushing the boundaries of VLM capabilities and boosting model performance in few-shot scenarios. However, these methods are all based on CLIP and do not investigate the performance of models on pathological VLMs. Moreover, due to the large number of patches in a WSI, they have to focus solely on the text and neglect visual prompt tuning. Additionally, they do not consider the crucial contextual information in WSI. Although ViLa-MIL takes into account multi-scale information, it merely integrates information using a late fusion approach without fully exploring the interactions between these scales.
We validated our proposed MSCPT on both general VLM (i.e., CLIP) and pathology-specific VLM (i.e., PLIP). By utilizing the zero-shot capability of VLM to initially select a subset of patches closely related to cancer, we then conducted visual prompt tuning on these patches. Additionally, we adopted an intermediate fusion approach to integrate pathology prior knowledge at multi-scale, leveraging the transformer layers to hierarchically learn the relationships between them. Ultimately, we also utilized image-text similarity to construct GNNs to capture contextual information within the WSI.
III Method
@@
In this section, we introduce our few-shot weakly-supervised WSI classification model, named Multi-scale and Context-focused Prompt Tuning (MSCPT), as illustrated in Fig.  2. MSCPT utilizes a dual-path asymmetric structure as its foundation while conducting hierarchical prompt tuning on both textual and visual modalities.
III-A Problem Formulation
@@
Given a dataset $X=\left\{X_{1},X_{2},...,X_{N}\right\}$ consisting of $N$ WSIs, each WSI is cropped into non-overlapping small patches, named instances. All instances belonging to the same WSI collectively form a bag. In weakly-supervised WSI tasks, only the labels of bags are known. The labels of the bags $Y_{i}\in\left\{0,1\right\},\ i=\{1,2,...N\}$ and the label of each instance $\{y_{i,j},j=1,2,\ldots M_{i}\}$ have the following relationship:
@@
$Y_{i}=\left\{\begin{array}[]{cc}0,&\quad\text{if}\sum_{j}y_{i,j}=0,\\ 1,&\quad\text{else}.\end{array}\right.$ (1)
III-B Review of CLIP and Patch Selection
III-B1 Review of CLIP

### v2

@@ Abstract
Multiple instance learning (MIL) has become a standard paradigm for the weakly supervised classification of whole slide images (WSIs). However, this paradigm relies on using a large number of labeled WSIs for training. The lack of training data and the presence of rare diseases pose significant challenges for these methods. Prompt tuning combined with pre-trained Vision-Language models (VLMs) is an effective solution to the Few-shot Weakly Supervised WSI Classification (FSWC) task. Nevertheless, applying prompt tuning methods designed for natural images to WSIs presents three significant challenges: 1) These methods fail to fully leverage the prior knowledge from the VLM’s text modality; 2) They overlook the essential multi-scale and contextual information in WSIs, leading to suboptimal results; and 3) They lack exploration of instance aggregation methods. To address these problems, we propose a Multi-Scale and Context-focused Prompt Tuning (MSCPT) method for FSWC task. Specifically, MSCPT employs the frozen large language model to generate pathological visual language prior knowledge at multiple scales, guiding hierarchical prompt tuning. Additionally, we design a graph prompt tuning module to learn essential contextual information within WSI, and finally, a non-parametric cross-guided instance aggregation module has been introduced to derive the WSI-level features. Extensive experiments, visualizations, and interpretability analyses were conducted on five datasets and three downstream tasks using three VLMs, demonstrating the strong performance of our MSCPT. All codes have been made publicly accessible at https://github.com/Hanminghao/MSCPT.
{IEEEkeywords}
@@ 1 Introduction
Developing automated analysis frameworks using Whole Slide Images (WSIs) is crucial in clinical practice [1, 2, 3], as WSIs are widely regarded as the “gold standard” for cancer diagnosis, typing, staging, and prognosis analysis [4]. Given the enormous size of WSIs (roughly 40,000 × 40,000 pixels), multiple instance learning (MIL) [5] has become the dominant method. As shown in Fig. 1a, traditional MIL-based methods typically follow a four-step paradigm: patch cutting, feature extraction, feature aggregation, and classification. Most MIL-based methods are conducted under weak supervision at the bag level, as creating instance-level labels is quite labor-intensive [6, 7]. This weak supervision paradigm has led to a problem: a large number of WSIs are required to train an effective model [8, 9]. In clinical practice, patient privacy concerns, rare diseases, and difficulties in preparing pathology slides make accumulating a large number of WSIs very challenging [10, 11].
Vision-Language models (VLMs) have shown excellent generalization ability to downstream tasks [12, 13, 14, 15, 16, 17]. Recently, researchers have proposed specialized VLMs for analyzing pathological images, including MI-Zero [18], PLIP [19], and Conch [20]. These VLMs, extensively pre-trained on abundant image-text pairs, contain significant prior knowledge. If the prior knowledge of VLMs can be fully exploited with a few training samples, it can partially alleviate the data scarcity problem in WSI classification tasks. Therefore, we aim to explore a novel “data-efficient” method based on the VLMs to improve the model’s performance on the Few-shot Weakly Supervised WSI Classification (FSWC) [6] task.
@@
However, existing prompt tuning methods for natural images (such as CoOp [22], CoCoOp [21], and MetaPrompt [23]) are only effective for single images (i.e., patch-level). Since each WSI typically contains tens of thousands of patches, these methods are ineffective for WSI-level tasks. Also, studies indicate that the multi-scale information [24] and the contextual information [25] in WSIs play a significant role in cancer analysis, but those methods fail to capture this crucial information. Additionally, in training VLMs, the image-text pairs contain more than just information about the category. They also include more details about the image, such as contextual properties of the object [12] and descriptions of the cellular microenvironment [19, 20]. However, existing prompt tuning methods have primarily focused on image category information without emphasizing a detailed image content analysis, which has left the full potential of pathological VLMs underexplored.
@@
To address the aforementioned issue, we propose Multi-Scale and Context-focused Prompt Tuning (MSCPT) for WSI classification in weakly supervised and few-shot scenarios. Our framework fully leverages the characteristic of VLM training with image-text pairs at dual magnification scales: 1) At low magnification, we provide the VLM with pathological visual descriptions at the tissue level (such as the infiltration between tumor tissue and other normal tissues); 2) At high magnification, pathological visual descriptions at the cellular level (such as cell morphology, nuclear changes, and the formation of various organelles) are provided to the VLM. These pathological visual descriptions at multiple scales can help VLM identify regions that are helpful for cancer analysis and achieve optimal results even with limited training samples.
@@
As illustrated in Fig. 1c, the core idea behind developing MSCPT is to incorporate prior knowledge at the tissue and cellular scales into the WSI-level tasks. Specifically, we first use a frozen large language model (LLM) to generate multi-scale pathological visual descriptions as prior knowledge.
@@
Secondly, we design a Multi-scale Hierarchical Prompt Tuning (MHPT) module to combine pathological visual descriptions from multiple scales hierarchically to enhance prompt effectiveness. Inspired by Metaprompt [23], a dual-path asymmetric framework is adopted, asymmetrically freezing the image encoder and text encoder at different scales for prompt tuning. This asymmetric framework enables us to freeze half of the encoder to reduce the number of trainable parameters. Specifically, MHPT contains low-level and high-level prompts for both low and high-magnification visual descriptions, as well as global trainable prompts. The MHPT module employs the transformer layers in the text encoder to effectively learn the interactions among three distinct prompts.
@@
Furthermore, the Image-text Similarity-based Graph Prompt Tuning (ISGPT) module is introduced to extract contextual information. Contextual analysis is crucial for WSI analysis [26, 1]. Researchers typically employ Graph Neural Networks (GNNs) or Transformer architectures to learn the context of WSIs. However, due to the large size of WSIs and the substantial computational resources required by the Transformer paradigm, these methods are not suitable for the FSWC task. Therefore, MSCPT adopts GNNs for lightweight contextual learning. Specifically, we do not follow previous approaches [26, 27] of using patch positions or patch feature similarity to construct GNNs. Instead, we propose to use the similarity between patches and pathological visual descriptions as the basis for building GNNs. We demonstrate that image-text pairs in GNNs more effectively capture global features than methods relying on patch positions and image similarity, as confirmed by corresponding ablation experiments.
@@
Finally, motivated by the powerful zero-shot capabilities of VLMs [19, 20, 18], we propose aggregating instances based on the similarity between image patches and pathological visual descriptions. The Non-Parametric Cross-Guided Pooling (NPCGP) module, utilizing the Top-K algorithm for instance aggregation, is introduced to further reduce the risk of overfitting in few-shot scenarios and endows MSCPT with human-readable interpretability. Overall, our contributions are:
@@
    MSCPT demonstrates that more detailed high-level concepts from pathological descriptions combined with low-level image representations can enhance few-shot weakly supervised WSI classification.
@@
    MSCPT achieves excellent performance by introducing only a limited number of trainable parameters ($0.4\%$ to $0.9\%$ of the pre-trained VLM). Additionally, MSCPT is applicable to fine-tune any VLMs for WSI-level tasks.
@@
    Extensive experiments, visualizations, and interpretability analyses conducted on five datasets spanning three downstream tasks and evaluated using three different VLMs confirm that MSCPT achieves state-of-the-art performance in few-shot scenarios, outperforming other traditional methods and prompt-tuning methods.
@@
Figure 2: We develop MSCPT based on the dual-path asymmetric framework, which inputs patches and pathological visual descriptions from multi-scale to different encoders. MSCPT utilizes a large language model to generate multi-scale pathological visual descriptions. These descriptions are combined using Multi-scale Hierarchical Prompt Tuning (MHPT) to integrate information across multiple scales. Then, Image-text similarity-based Graph Prompt Tuning (ISGPT) is employed to learn context information at each scale. Finally, Non-Parametric Cross-Guided Pooling (NPCGP) aggregates instances guided by pathological visual descriptions to achieve the final Whole Slide Image classification result.
2 Related Work
2.1 Multiple Instance Learning in Whole Slide Images
@@
Due to the high resolution of Whole Slide Images (WSIs) and the challenges of detailed labeling, weakly supervised methods based on Multiple Instance Learning (MIL) have emerged as the mainstream for WSI analysis. The MIL-based methods treat a WSI as a bag and all patches as instances, considering a bag positive if it contains at least one positive instance. Within the MIL framework, a step is required to aggregate all instances into bag features. The most primitive aggregation methods are non-parametric mean pooling and max pooling. However, since disease-related instances are a small fraction [28], those non-parametric aggregation methods treated all instances equally, causing useful information to be overwhelmed by irrelevant data. Subsequently, some attention-based methods (such as ABMIL [29], DSMIL [30], and CLAM [28]) were introduced, assigning different weights to each instance and aggregating them based on the weights. Furthermore, MIL methods based on Graph Neural Networks (GNNs) [26, 27] and Transformers [1, 31] have also been proposed to capture both local and global contextual information of WSIs. Those methods have shown significant improvements in recent years. Still, the cost of enhancing model performance is the increase in parameters, requiring a large amount of data to train a well-performing model. In many cases, training data faces a scarcity issue. Therefore, this paper proposes MSCPT, which leverages Vision-Language models combined with pathological descriptions from Large language models to enhance the performance in few-shot scenarios.
2.2 Vision-Language Models
@@
Vision-Language models (VLMs) are rapidly developing in various fields. During training, VLMs use contrastive learning to reduce distances between paired image-text pairs and increase distances between unpaired ones. The creators of CLIP [12] collected over 400M image-text pairs from the internet and used contrastive learning to align them, resulting in compatibility across various tasks. Compared to natural images, gathering pairs of pathological images and corresponding descriptions is challenging. To address this issue, MI-Zero [18] first pre-trained image and text encoders using unpaired data and then aligns them in a common space using 33,480 pathological image-text pairs. Huang et al. gathered over 208K pathological image-text pairs from Twitter and LAION [32] and developed PLIP [19]. Lu et al. trained Conch [20] on over 1.17M pathological image-text pairs, and this model performs well on downstream tasks. Pre-trained VLMs have significant potential, but effective methods to leverage them for WSI-level tasks are lacking. In this paper, we propose using pathological visual descriptions as prior knowledge to unleash the potential of VLMs.
2.3 Prompt Tuning in Vision-Language Models
@@
Prompt tuning has demonstrated remarkable efficiency and effectiveness, whether in text or multimodal domains [22, 21, 14]. CLIP demonstrated remarkable zero-shot performance with hand-crafted prompts, but the results can vary significantly depending on the prompt used due to their sensitivity to changes. Therefore, CoOp [22] and CoCoOp [21] proposed that the model itself should determine the choice of prompts. Khattak et al. argued that optimizing prompt tuning within a single branch is not ideal. They introduced MaPLe [14] to enhance the alignment between visual and language representations. However, these innovative methods are highly applicable to natural images but do not consider the enormous size of WSIs and the crucial multi-scale and contextual information needed for WSI analysis. To the best of our knowledge, TOP [6] was the first study to explore fine-tuning CLIP for FSWC tasks. Shortly thereafter, Shi et al. introduced ViLa-MIL [33], which builds upon CLIP and enhances WSI classification by incorporating multi-scale language prior knowledge. These pioneering studies have significantly advanced the utility of VLMs and improved model performance in few-shot WSI classification scenarios.
However, both TOP and ViLa-MIL are CLIP-based and do not evaluate the performance of models on pathological VLMs. Moreover, due to the large number of patches in a WSI, these approaches primarily emphasize text while overlooking visual prompt tuning and the crucial contextual information within the WSI. Although TOP focuses on instance-level and batch-level prompt tuning, it does not incorporate multi-scale information. ViLa-MIL, on the other hand, incorporates multi-scale information but employs a late fusion strategy, which limits its ability to fully explore interactions across scales. Furthermore, while ViLa-MIL introduces category-specific priors, it is constrained by the context length of the CLIP text encoder (77 tokens), restricting the amount of prior knowledge it can utilize. Recent advances in pathology-specific VLMs [20, 34, 35, 36] have significantly enhanced their ability to capture domain-specific pathological knowledge, motivating the development of MSCPT to better leverage detailed and effective prior knowledge.
3 Method
@@
In this section, we introduce our few-shot weakly supervised WSI classification model, named Multi-Scale and Context-focused Prompt Tuning (MSCPT), as illustrated in Fig.  2. MSCPT utilizes a dual-path asymmetric structure as its foundation while conducting hierarchical prompt tuning on both textual and visual modalities.
3.1 Problem Formulation
@@
Given a dataset $X=\left\{X_{1},X_{2},...,X_{N}\right\}$ consisting of $N$ WSIs, each WSI is cropped into non-overlapping small patches, named instances. All instances belonging to the same WSI collectively form a bag. In weakly supervised WSI tasks, only the labels of bags are known. The labels of the bags $Y_{i}\in\left\{0,1\right\},\ i=\{1,2,...N\}$ and the label of each instance $\{y_{i,j},j=1,2,\ldots M_{i}\}$ have the following relationship:
@@
$\small Y_{i}=\left\{\begin{array}[]{cc}0,&\quad\text{if}\sum_{j}y_{i,j}=0,\\ 1,&\quad\text{else}.\end{array}\right.$ (1)
3.2 Review of CLIP and Patch Selection
3.2.1 Review of CLIP